OpenStudy (anonymous):

Simplify the following:

OpenStudy (anonymous):

\[\frac{ 3a }{ a+3 } - \frac{ (a-4) }{ a-2 }\]

OpenStudy (anonymous):

Common denominator is ....?

OpenStudy (anonymous):

Ahhhhh dang...you're right.

OpenStudy (anonymous):

Hey, I'm waiting to see your multiplication ;)

OpenStudy (anonymous):

It didn't work so I have no idea.

OpenStudy (anonymous):

I thought of making both denominators equal but I can't x_x

OpenStudy (anonymous):

Show me so I can know where to fix it?

OpenStudy (anonymous):

\[\frac{ 6a }{ 2a + 6 } - \frac{ -3a+ 12 }{ 3a + 6 }\]

OpenStudy (anonymous):

Ahhh!!! sorry, on the right, they're negatives.

OpenStudy (anonymous):

Common denominator is product of (a +3) ( a-2): -> 3a ( a -2 ) = .... -> (a - 4) ( a +3) = ...

OpenStudy (anonymous):

Just follow me, work step by step fill out the blank!

OpenStudy (anonymous):

Haha, I'll try.

OpenStudy (anonymous):

Please post here!

OpenStudy (anonymous):

Post what?

OpenStudy (anonymous):

-> 3a ( a -2 ) = .... ?

OpenStudy (anonymous):

\[3a ^{2} - 6a\]

OpenStudy (anonymous):

So it's just cross multiply?

OpenStudy (anonymous):

Great :) -> (a - 4) ( a +3) = ... ?

OpenStudy (anonymous):

Exactly right :)

OpenStudy (anonymous):

\[a ^{2} -12\]

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

I forgot some!!!

OpenStudy (anonymous):

FOIL method!

OpenStudy (anonymous):

\[a ^{2} -4a + 3a\] -12

OpenStudy (anonymous):

-> (a - 4) ( a +3) = a² -4a + 3a - 12 = ...

OpenStudy (anonymous):

Can you simplify -4a + 3a = ....?

OpenStudy (anonymous):

-a

OpenStudy (anonymous):

So the answer is....

OpenStudy (anonymous):

-> (a - 4) ( a +3) = a² -a - 12

OpenStudy (anonymous):

\[\frac{ 2a ^{2} -7a -12 }{ (a+3) (a-2) }\]

OpenStudy (anonymous):

Now after cross product, subtract the numerators: 3a²- 6a - (a² -a - 12) = 2a² -6a +a +12 = ...

OpenStudy (anonymous):

So I'm right?

OpenStudy (anonymous):

-6a + a = ....?

OpenStudy (anonymous):

Well the answer on the problem says 7a

OpenStudy (anonymous):

I don't see how they get 7a, probably typo :)

OpenStudy (anonymous):

Maybe.

OpenStudy (anonymous):

I have 3 more and I tried one this way and it doesn't work.

OpenStudy (anonymous):

In short, learn how to subtract unlike fraction by cross product!

OpenStudy (anonymous):

Open the new posts for new questions!

OpenStudy (anonymous):

Make sure your post is correct!